OpenStudy (anonymous):

Which equation represents a line that intersects the line 2y − 4x = 3? -2y + 4x = 1 y − 3x = 2.5 y − 2x = 3 -2y − 4x = 3

OpenStudy (gorv):

from equation give find the value of either x or y and put in givrn equation one by one if they cut their will be some value of x and y

OpenStudy (igreen):

You can graph it here: https://www.desmos.com/calculator Both B and D intersect the line.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @rainelisl

OpenStudy (anonymous):

Find the intersection the best way would be by elimination in such a scenario multiply equation 1 by 2 to get the 4x so we can subtract so 4x+2y+2=0 4x+y-3=0 subtract equations y+5=0 y=-5 plug in any equation 4x-5-3=0 4x-8=0 4x=8 x=2 the solution set is (2,-5)